Summer Service 2005

Diary of my summer 2005 equipment service


Once it reaches June in my latitude serious observation becomes very difficult. True astronomical darkness does not return until late July and the fleeting hours of twilight are clustered around the wee small hours. This makes it a perfect time to take the equipment down from the observtory and give everything a good service to maximise observations when darkness returns.

Now follows some chronological notes of my doings:

31st May 2005

Friends visiting this evening, but I decided to make a start by unplugging everything in the observatory, wrapping the cameras and storing them away. The OTA was moved inside the house to an area prepared in the workshop. I removed the primary mirror cell from the back of the scope, and removed the mirror from the cell. The cell on my cheap newtonian is dreadful - perhaps I can do something to improve this and get rid of the nasty retaining clips. Inspecting the mirror I find its really very murky. Lots of dried on gunk and dust. Cleaning that will be the last job! The mirror was stored in a very safe place.

1st June 2005

I decided to start on the focuser area. Last summer I cleaned up the focus drawtube and applied some decent grease. However, I've noticed a bit of image shift creeping in and the motor struggling at times. The motor gears are a bit rusty, and there are a lot of wires "floating" about up near the focuser. All needs tidying up. s
But first that drawtube. I fixed it up in the drill press and ran over it with the 120/400/600/800/1200 grade sandpaper and then some metal polish. Much smoother looking! Not perfect by a long way, but probably good enough for my purposes.

8th June 2005

A weeks break from any work on the scope. Not good! Only a couple of weeks until solstice - I was hoping to be further on than this, but friends will have weddings in the summer! Tonight I was able to strip the focuser down and, with a little sandpapering here and there come up with an arrangement of PTFE shims that gives as close to a zero image shift focuser as I can manage with a good old fashioned drawtube focuser. I cannot detect any wobble, and neither can the spirit level. Next job is to clean up the focuser in white spirit to get off all the aluminium fillings and last years grease. Hopefully next session the focuser will be done and we can turn to other matters.

9th June 2005

Moving forwards once more. The focuser is cleaned and regreased. All the screws are cleaned up and dampened with WD40. Its all back together, with the PTFE glued into place and the whole thing roughly adjusted. Focuser is now wrapped in cling film and put in the drawer until later.
I've also removed the front ring of the scope and the spider, and all its wiring for the dew heaters. The secondary has joined the primary in the safe drawer. The inside of the scope is rather grotty. I've taken it out of the rings and started experimenting with various paints to see what gives me a good finish in the inside as a sort of flocking. I'm not convinced about flocking paper etc - its very hard to glue such things into the scope tube, and any material is going to drop fibres into the OTA.
The dec axis has been removed from the mount and cleaned up. I'm debating removing the RA axis - it seems to be working quite well as it is, and I don't want to upset it. The next project is to devise a way of bolting the dovetail bar, saddleplate and DEC axis into one mechanical piece for rigidity. There is a serious crack in one of the screwholes for the saddle plate in the top of the dec axis, and I want something more reliable.

10th June 2005

Hmm, change of mind on the flocking. I have located some synthetic flocking material. Lets see what its like when it turns up. I've also drilled some holes in the saddle plate (eeek!) to add some bolts to give additional strength. Most of the evening has been spent playing around with some bits of aluminum to build a rough mirror cell. Worked out quite well, execpt I got some of my geometry wrong, and need to make some adjustments. However, pleasing progress.

17th June 2005

Frantic week. I have ordered and received some flocking material for the inside of the OTA. To enhance the scruffy looking outside I have decided to paint it white. This will make things more visible in the dark. I've sanded down the existing black gloss and applied a tin of grey primer. I will leave that tonight and put some white on it in the morning. Mother is visiting tomorrow - but whilst she is doing the gardening thing I am going to concentrate on the observatory and getting the mount back together. Sanded down the outside of the OTA with 400 and 800 paper, and gave it good coat of primer.

18th June 2005

Very busy today. Sanded down the primer with 800 and 1200 paper, and then sprayed it white. Looks very nice ;-) Mother then helped me glue on the flocking material with some "moveable" impact adhesive. Seems to have stuck on, not too much of a mess. Spent the evening finishing dismantling the mount and cleaning it all in white spirit.

19th June 2005

Very hot in the UK today. The inside of the observatory has hit 40C. I've repolished the bearing surfaces of the mount, regreased and reassembled. Spent the usual hour messing around with the worm gear adjustment. Too hot to put the mount in the obs today. I've also been building a mirror cell. Found some large springs for the collimation bolts and I *think* what I have will work. Its a very crude 6 point cell - but ANYTHING will be an improvement on the existing cell. This supports the mirror in three points on its circumference. It also has the dreaded "mirror clips". To make matters worse, collimation involved lossening one nut, adjusting another and then tightening the first. Very fiddly. My new cell will hopefully improve on all these matters.

20th June 2005

Finalised construction of the mirror cell. All looking good. Beginning to experiment with methods of attaching the mirror to the cell.

21st June 2005

Summer Solstice! A suitable task tonight - started to put everything back together. Concentrated on re-attaching the focuser assembly. Careful use of a laser and a ruler got the focuser pretty square to the OTA.

22nd June 2005

Ripped the camera appart, ready for reconstruction with new case design. Hacksawed up a large plate of copper that may or may not be used.

23rd June 2005

Drilled some holes in the mirror cell to allow sideways support/adjustment screws. Mounted copper plate in camera box.

24th June 2005

Moment of Plop! Glued mirror to mirror cell. Used silcone glue.

27th June 2005

After several days leaving the mirror cell to dry, I gave it a poke this evening, it seems to be quite firmly set. Cleaned the mirror.

28th June 2005

Rebuilt the telescope! Put the mirror cell back together with the OTA and mounted the whole thing on the EQ head on a tripod. EQ RA Gear far too tight :( Glued and screwed the finderscope bracket back on to the OTA.

30th June 2005

Adjusted out the stiff RA gear - just a smidge of backlash, but autostar should be able to deal with that. Did a better collimation. Started doing the wiring at the top of the scope. Halfway throuh this I stopped and took the scope outside with some friends to watch Venus, Mercury and Jupiter. Saturn proved invisible. First light images looked OK. When and looked at Arcturus. Nothing too scarey in the optics department. Hopefully we will be back in the observatory this weekend!

3rd July 2005

Going to the observatory did not happen. Appart from live8, not much else happened this weekend until Sunday night when I made good progress reconstrucing the camera and waterproofing the cold chamber. Also repaired some damage to the camera. Hopefully we will get the camera working tonight.

8th July 2005

Telescope now installed in observatory. Picture.

9th July 2005

First proper night out with the scope. Concentrated on trying to get everything working in concert. Main problem was focusing the guidescope and leveling of CCD chip.

10th July 2005

Sorted out the CCD chip and figured out how to focus the guidescope using a terrestial target. Rigged up the camera in the observatory with the scope closed and tested the darkframes.

5x10s frames warm camera. Histogram stretched X4 in photoshop. (ie 0-64 range) Image

5x10s frames cold camera. Same as above. Histogram stretched X4 in photoshop. (ie 0-64 range) Image


Well, thankfully thats all over with. Turned out to be a lot more effort than I had first planned. Of course, I won't really know if everything is working as well as I hope until we get some decent conditions.